OpenStudy (anonymous):

Solve for x. x^2/3=9

OpenStudy (anonymous):

x= 3(sqrt(3))

OpenStudy (anonymous):

can i assume you have \[x^{2/3}=9\]

OpenStudy (anonymous):

I don't think he means that.

OpenStudy (anonymous):

If so raise both sides to the 3/2 power\[\left( x^{2/3} \right)^{3/2}=9^{3/2}\]\[x=\left( 9^{1/2} \right)^3\]

OpenStudy (anonymous):

\[x=\left( \sqrt{9} \right)^3=3^3=27\]

OpenStudy (anonymous):

! :) thank you

OpenStudy (anonymous):

:})